OpenStudy (anonymous):

WIL GIVE MEDAL! HELP Quadrilateral ABCD has coordinates A (3, -5), B (5, -2), C (10, -4), D (8, -7). Quadrilateral ABCD is a (4 points)

OpenStudy (anonymous):

rectangle, because opposite sides are congruent and adjacent sides are perpendicular square, because all four sides are congruent and adjacent sides are perpendicular parallelogram, because opposite sides are congruent and adjacent sides are not perpendicular rhombus, because all four sides are congruent and adjacent sides are not perpendicular

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

Solution 1>> graph it. Solution 2>> find the distance between the points. little busy :)) try them u'll get answer

OpenStudy (anonymous):

oksy can you help me with one more please?@princeharryyy

OpenStudy (anonymous):

@eninone can you help me with another one?

OpenStudy (anonymous):

@seth23 whynot, what's problem? post it.

OpenStudy (anonymous):

Find the midpoint of diagonal BD. https://adomvirtual.brainhoney.com/Resource/27440840,0/Assets/74658_53c6c673/0601_g1_q1.jpg

OpenStudy (anonymous):

@eninone

OpenStudy (anonymous):

(1, 3.5) (1.5, 4) (2, 3.5) (2, 4)

OpenStudy (anonymous):

\[cordinate \ of \ point \ B \ are \ (0,6) \ and \ point \ D\ are\ (3,2)\\ now \ using\ distance\ formula\ btw \ B \ and \ D\\ BD=|\sqrt{(0-3)^2+(6-2)^2}|\\ \implies \ BD= |\sqrt{3^2+4^2}|\\BD=|\sqrt{9+16}|\\BD=|\sqrt{25}|\\BD=5\ unit\\ \let \ \mid \ points\ of \ BD \ be\ (x,y)\\ then\ x=\frac{ 0+3 }{ 2}= \frac{ 3 }{ 2 }=1.5\\ y= \frac{ (6+2) }{ 2}=\frac{ 8 }{ 2 }=4\\ \therefore\ midpoints\ of\ BD \ are \ (1.5\,\ 4)\ Q.E.D.\\\\ -----LOGICFALL\ Co.-----\]

OpenStudy (anonymous):

oh! no, no need to find distance BD, mistakely done! Just see the last part. @seth23 ANS= (1.5, 4)

OpenStudy (princeharryyy):

Sorry ! I was on call.